OpenStudy (anonymous):

Prove that the series [ln((n+1)/n)] from n=1 to infinity is divergent using the fact that ln((n+1)/n) = ln(n+1)-ln(n) and write out a few terms.

myininaya (myininaya):

Did you try to write out some terms?

OpenStudy (anonymous):

Yup. I got {0.7, 0.4, 0.4, 0.2...}

myininaya (myininaya):

I recommend not simplifying...

myininaya (myininaya):

like use the hint given

OpenStudy (anonymous):

Should I use one of the tests? I was just thrown off by the info given.

myininaya (myininaya):

write out the first few terms of the series

myininaya (myininaya):

\[\lim_{n \rightarrow \infty } \sum_{i=1}^{n} (\ln(n+1)-\ln(n))\] begin by replacing n with 1 then n with 2 then n with 3 then n with 4 ... You should see a pattern Do not use a calculator

OpenStudy (anonymous):

It looks like it's converging toward zero. Is that off-base?

myininaya (myininaya):

All I'm asking you to do is do this: ln(1+1)-ln(1) + ln(2+1)-ln(2) + ln(3+1)-ln(3) + ln(4+1)-ln(4) + ln(5+1)-ln(5) + ln(6+1)-ln(6) what do you notice?

OpenStudy (anonymous):

Oh! I see that both parts are increasing!

myininaya (myininaya):

:( that isn't what I'm trying to get you to see

myininaya (myininaya):

look ln(2)-ln(2)=? ln(3)-ln(3)=? ln(4)-ln(4)=?

OpenStudy (anonymous):

Oh! I didn't look at it as a whole. It's all actually adding up to zero!

myininaya (myininaya):

well except for one term

myininaya (myininaya):

what term is left over?

OpenStudy (anonymous):

The final term in the sequence, right? Ln(infinity)? :/

myininaya (myininaya):

do you think ln(n+1) or ln(n) is what is left over?

myininaya (myininaya):